Calibration progress

Speaker : Gene V. Buren ( BNL )


Talk time : 12:40, Duration : 00:20

Calibration activities and status updates since Calibration status for the multiple years productions ... highlighted in red...


Recent R&D on which calibrations have dependence:

  1. Inclusion of HFT material accounting in Sti
  2. Static TPC distortion/alignment study
  3. Asymmetric TPC SpaceCharge & GridLeak
    • Asymmetric colliding systems, e.g. dAu, CuAu
    • You do not have access to view this node, e.g. HFT support structures on east-side only

A menu of datasets:

Run 13 pp510 (second half)
  • TPC SpaceCharge & GridLeak done (including confirmation pass)
  • BeamLine calibration done
  • TPC dE/dx calibration checked
    • Existing calibration from first half of dataset checked with second half and declared sufficient
  • TPC hit errors calibrated
    • Still to be uploaded to database
  • TOF calibration
    • Production done using trigger-filter for speeding up production, as was done for first half of dataset
    • VPD calibration done
    • BTOF alignment ongoing
  • All calibrations completed and ready for production next week
Run 12 UU193
  • Testbed for new TPC alignment (symmetric system)
  • Will require the full suite of calibrations re-done
  • To vet alignment, only re-doing a few:
    • TPC T0 done
    • TPC SpaceCharge & GridLeak done
    • TPC Drift Velocity will finish this week
      • Added to the to-do list since last time as east-west differences could be relevant and calibration is (supposed to be) simple
  • Imminent: test production for h-/h+ as a metric for TPC alignment performance
Run 14 AuAu15
  • Preview production in May had You do not have access to view this node, old TPC alignment, no HFT material accounting
  • TPC T0 re-calibration done without new alignment
  • New alignment and new material accounting will however require full suite of calibrations re-done
    • TPC SpaceCharge & GridLeak was not large (low collision energies had low luminosity), so asymmetric SC & GL not critical
  • TOF group is exploring using this data to calibrate for Run 14
Run 14 AuAu200
  • Dependencies on all above R&D
  • Full suite of final calibrations yet to be started
Run 14 HeAu200
  • Similar to Run 14 AuAu200
Run 12 CuAu200
  • Awaiting R&D on asymmetric TPC SpaceCharge & GridLeak
  • TOF, TPC dE/dx, TPC hit errors to follow
Run 8 dAu200
  • Similar to Run 12 CuAu200

Misc related upcoming R&D (pieces of the big puzzle: we have not forgotten these things, but priorities must be kept)

  • TPC calibration improvements using GMT
    • Investigating observations of ~2 cm split between east-TPC tracks and west-TPC tracks in GMT
  • TPC calibration improvements using HFT
    • You do not have access to view this node already started
    • Dependency: tracking with HFT hits


-Gene